OpenStudy (anonymous):

I have homework about Simplifying Rational Expressions I will fan and medal 😝✌️

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

http://prntscr.com/9okjrd

OpenStudy (anonymous):

okay hold up i would have to use my other computer to login.....because the one am with right now is blocked

OpenStudy (anonymous):

ok :P

OpenStudy (anonymous):

okay lets do this

OpenStudy (anonymous):

WB 😜

OpenStudy (anonymous):

look closely to the question what is your first step

OpenStudy (anonymous):

Factor both the numerator and denominator

OpenStudy (anonymous):

lets ignore the alphabet first and concentrate on the number only

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

now factorize the number to its lowest form

OpenStudy (anonymous):

would the answer be \[\frac{ 2 }{3}\]

OpenStudy (anonymous):

alright since u've done that spread out the alphabet close to this number

OpenStudy (anonymous):

okay wait lets cut the long method short

OpenStudy (anonymous):

will you like the long method or short

OpenStudy (anonymous):

ok 😐

OpenStudy (anonymous):

if your answer is 2z^4/3z^4 and you are told that its not yet done what will you do next

OpenStudy (anonymous):

simplify

OpenStudy (anonymous):

now what are you canceling out then

OpenStudy (anonymous):

let me give you a hint: 4ab/5ab=4/5 6a^2/11a^2=6/11

OpenStudy (anonymous):

ummm

OpenStudy (anonymous):

i think i got it from here

OpenStudy (anonymous):

okay hope the hint helped

OpenStudy (anonymous):

it did thx 😆

OpenStudy (anonymous):

so what is your final answer

OpenStudy (anonymous):

im still working on it